The average molar mass of components of air mainly diatomic gas and diatomic oxygen gas is about 29g/mol what is the volume of 1kg of air at atmospheric pressure and 20C


1
Expert's answer
2020-06-08T10:28:33-0400

According to the Ideal Gas Law, the volume of one mole of gas is


"V=\frac{nRT}{p}=\frac{1\cdot8.314\cdot(273+20)}{101325}=24\cdot10^{-3}\text{ m}^3."

Therefore, the volume of gas does not depend on its molar mass.